# RefDrone: A Challenging Benchmark for Referring Expression Comprehension in Drone Scenes

### v1

@@ 3.4 Dataset comparison
Table 1 compares RefDrone with existing REC datasets. RefDrone stands out by its average number of referred objects and its use of LMM for expression generation. The expressions offer richer contextual details compared to template-based or human-annotated expressions. While RIS-CQ [13] employs LLM for expression generation, it lacks visual content during the generation process. This often results in expressions that, despite being linguistically complex, may be disconnected from visual content. Similarly, RSVG [47], although it focuses on small targets, falls short in terms of expression quality. In contrast, RefDrone comprehensively presents three challenges, establishing it as a challenging benchmark in REC tasks.
@@ 3.5 Evaluation metrics
We introduce instance-level metrics extending traditional REC metrics to address multi-target challenges. Previous benchmarks primarily focus on image-level metrics, which are sufficient for single-target or few-target samples. However, these metrics fail to handle expressions referring to multiple target objects—potentially scaling up to 100 instances per expression in our task. To address this limitation, we introduce instance-level metrics that provide a more granular evaluation of a model’s capability to identify individual target objects.
Instance-level metrics: Accinst. and F1inst. measure the accuracy of individual bounding box predictions. We compute the intersection over union (IoU) between the true bounding boxes and the predicted bounding boxes. An IoU $\geq$ 0.5 is considered a true positive (TP), otherwise it is a false positive (FP). Unmatched true bounding boxes are counted as false negatives (FN). For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP). We calculate:
$\text{Acc}_{inst.}=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
and
@@
Image-level metrics: Accimg. and F1img. evaluate the overall accuracy of predictions per image. An image is considered a true positive (TP) if all predictions match the true bounding boxes, otherwise it is a false positive (FP). For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP). We calculate Accimg. and F1img. using the same formulas as Accinst. and F1inst., respectively.
4 Proposed method
We introduce Number GroundingDINO (NGDINO) to address multi-target and no-target challenges based on GDINO [28]. Our insight is that utilizing number information of referred objects enhances model performance on these challenges. As illustrated in Figure 6, we introduce three components: (1) a number prediction head, (2) number-queries, and (3) a number cross-attention module.
@@ 4 Proposed method
4.1 NGDINO
Following GDINO, our model employs a dual-encoder-single-decoder architecture. The framework consists of an image backbone for visual feature extraction, a text backbone for text feature extraction, a feature enhancer for cross-modal fusion, a language-guided query selection module for detection-query initialization, and a cross-modality decoder for box refinement. Within decoder layers, detection-queries are fed into a self-attention layer, an image cross-attention layer to combine image features, a text cross-attention layer to combine text features, and an FFN layer. We mainly improve the decoder part, highlighted by the yellow box in Figure 6.
The number prediction head, structured similarly to the detection head with an FFN layer, predicts the number of referred objects from detection-queries. These detection-queries are used to predict the objects, containing number information. The number-queries are learnable embeddings designed to capture various numerical patterns. These queries are initialized randomly with dimensions (bs, length_nquery, ndim), where bs is the batch size, length_nquery represents the number of query lengths, and ndim denotes the feature dimension. The predicted number guides the selection of number-queries through the query selection in Algorithm 1. The selected number-queries have a fixed length of length_snquery.
We implement number cross-attention in parallel with self-attention to integrate the number information. In the number cross-attention, the selected number-queries serve as keys and values, while detection-queries serve as queries. The outputs are added to the self-attention features.
4.2 Loss function
[…]
""" Input: detection_queries:(bs,length_dquery,ndim) number_queries:(bs,length_nquery,ndim) Output: selected_number_queries:(bs,length_snquery,ndim) """ number_feat = FFN(detection_queries) # (bs, length_dquery, 1) number_feat = number_feat.mean(dim=1) # (bs, 1) number = sigmoid(number_feat) # (bs, 1) index = slice(length_snquery*number, length_snquery*(number+1)) selected_number_queries = number_queries[index] # (bs, length_snquery, ndim)
Algorithm 1 Query selection.
@@ 5 Experiments
5.1 Construction of RDAgent and NGDINO
RDAgent. A straightforward method is to adapt RDAgent for REC tasks. This adaptation automates the process by replacing the object input with predictions from an object detector and removing the human verification step. For a fair comparison, we employ Faster R-CNN [34] as the object detector, which is trained on the VisDrone2019-DET [49] dataset. The agent component utilizes GPT-4o [^1] https://openai.com/index/hello-gpt-4o/.
NGDINO. The NGDINO adopts a two-stage procedure to maintain training stability. First, we pre-train the number prediction head on the RefDrone dataset while initializing other components with GDINO [28] parameters. Then, we fine-tune the entire model.
5.2 Implementation details
We establish the benchmark with 13 representative models that can perform REC tasks, comprising 3 specialist models and 10 LMMs. The specialist models include MDETR [14], GLIP [19], and GDINO [28]. For LMMs, we evaluate Shikra [5], ONE-PEACE [38], SPHINX-v2 [21], MiniGPT-v2 [3], Ferret [44], Kosmos-2 [31], Griffon [48], Qwen-VL [1], CogVLM [41], and LLaVA [24]. Detailed model specifications are provided in the Supplementary Material.
Zero-shot evaluation details. For zero-shot evaluation, we use the original model checkpoints as provided in their respective papers. The implementations and checkpoints for GLIP and GDINO are obtained from the MMDetection [4].
Fine-tuning evaluation details. Our fine-tuning protocol maintains consistency across all experiments to ensure a fair comparison. We preserve the original learning strategies while excluding random crop augmentation due to its negative effect on position-sensitive samples. For LMMs, we employ the LoRA [11] fine-tuning strategy and follow the instruction tuning data structures. All fine-tuning experiments run for 5 epochs using 4 NVIDIA A100 GPUs.
@@ 5.3 Experimental results
Table 2: Experimental results of zero-shot baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks. GDINO-T and GDINO-B denote GroundingDINO [28] with Swin-Tiny and Swin-Base backbones. The best results in each group are denoted with bold.
Zero-shot results. Table 2 presents the zero-shot evaluation results, assessing the models’ domain generalization capability. CogVLM [41] demonstrates state-of-the-art performance across multiple metrics. However, several advanced models, including Shikra [5], ONE-PEACE [38], and SPHINX-v2 [21], are limited to outputting only a single bounding box, due to constraints in their pre-training data or output strategy. This restriction significantly impacts their performance in multi-target scenarios.
Table 3: Experimental results of fine-tuning baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks.
Methods F1inst. Accinst. F1img. Accimg. MDETR [14] 32.60 19.49 19.17 10.81 GLIP [19] 24.23 14.86 16.92 13.29 GDINO-T [28] 30.50 19.02 29.65 21.07 NGDINO-T (Ours) 33.34 20.98 32.45 22.84 GDINO-B [28] 31.96 19.99 31.69 22.26 NGDINO-B (Ours) 34.44 21.76 34.01 23.89 MiniGPT-v2 [3] 4.97 2.74 13.56 8.97 LLaVA [24] 6.00 3.63 14.43 11.57 Qwen-VL [1] 14.14 7.61 20.10 11.17 RDAgent (Ours) 58.14 41.13 37.07 23.54
Fine-tuning results. Table 3 illustrates the fine-tuning performance across different methods. The specialist model MDETR [14] shows strong performance in instance-level metrics, achieving comparable results to GDINO-B [28], but struggles in image-level understanding. Conversely, LMMs like Qwen-VL [1] demonstrate superior image-level comprehension (F1img.: 20.10%, Accimg.: 11.17%) while struggling with instance-level tasks (F1inst.: 14.14%, Accinst.: 7.61%). This performance disparity can be attributed to LMMs’ strong global image understanding capabilities but limited effectiveness in detecting small objects due to input resolution constraints.
The proposed RDAgent achieves superior results compared to existing approaches, particularly in instance-level metrics, surpassing GDINO-B [28] by 26.18% in F1inst. and 21.14% in Accinst.. The relatively smaller gains in image-level metrics (5.38% in F1img. and 1.28% in Accimg.) highlight the importance of instance-level metrics. However, RDAgent’s multi-step pipeline and reliance on ChatGPT responses result in longer processing time compared to end-to-end approaches.
The proposed end-to-end method, NGDINO, consistently improves over the baseline GDINO [28] across both backbone architectures. Figure 7 shows the visualization results comparing NGDINO and baseline GDINO, demonstrating NGDINO’s effectiveness with multi-target samples.
@@
Methods testA testB
Pr@0.5 $\uparrow$ N-acc. $\uparrow$ Pr@0.5 $\uparrow$ N-acc. $\uparrow$
MDETR* [14] 50.0 34.5 36.5 31.0
UNINEXT* [43] 46.4 49.3 42.9 48.2
GDINO-T [28] 45.69 79.02 44.83 76.69
NGDINO-T 46.05 83.17 45.57 78.08
To further validate the effectiveness of the proposed NGDINO, we conduct additional experiments on the gRefCOCO [22] dataset, which also includes multi-target and no-target samples (Table 4). NGDINO-T outperforms the baseline method GDINO-T, particularly in N-acc. metrics (by 4.15% in test A and 1.39% in test B), which evaluate accuracy on no-target samples. The limited improvements in Pr@0.5 metrics can be attributed to the relatively simple nature of multi-target samples in gRefCOCO, which typically contain only two objects per sample.
Figure 7: Visualization results comparing NGDINO and GDINO [28], where red boxes indicate true positives and yellow boxes denote false negatives.
Figure 8: Failure cases of NGDINO on RefDrone dataset. Red, green, and yellow boxes indicate true positives, false positives, and false negatives, respectively.
Table 5: Ablation study on RefDrone dataset for the effectiveness of each component in NGDINO, with Swin-Tiny backbone. NPH denotes number prediction head. NCA denotes number cross-attention.
NPH NCA Acc_$\_inst.$ Acc_$\_img.$ FPS 19.02 21.07 13.5 19.09 (0.07$\uparrow$) 21.18 (0.11$\uparrow$) 12.8 (0.7$\downarrow$) 19.51 (0.49$\uparrow$) 21.71 (0.53$\uparrow$) 12.9 (0.6$\downarrow$) 20.98 (1.96$\uparrow$) 22.84 (1.77$\uparrow$) 12.3 (1.2$\downarrow$)
Table 6: Ablation study for the impact of varying selected number query length. Params indicates additional parameters introduced.
@@
Table 7: Ablation study for the effectiveness of RDAgent.
Methods F1inst. Accinst. F1img. Accimg. ReCLIP [36] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAgent 58.14 41.13 37.07 23.54
@@ 5.4 Ablation studies
Ablations of NGDINO components. Table 5 presents the analysis of each component in NGDINO. The number prediction head alone shows a slight increase in Accinst. from 19.02% to 19.09%, indicating minimal impact of this auxiliary head on the referring tasks. When introducing the number cross-attention without number selection, we observe a more substantial improvement, with Accinst. increasing from 19.02% to 19.51%. This improvement is attributed to the additional parameters introduced in the decoder. The significant improvement is achieved through the combination of both the number prediction task and number cross-attention components. This contribution increases Accinst. from 19.02% to 20.98%. While the additional computational cost results in a minor decrease in inference speed (FPS), this trade-off yields a 1.96% increase in Accinst..
Ablations of query length. Table 6 analyzes the impact of varying the selected number query length. Utilizing a minimal query length of 1 lacks the capacity to capture complex numerical information. Conversely, extending the query length to 100 increases parameter count and introduces optimization difficulties, thereby adversely affecting performance. Through these experiments, we determine that a query length of 10 provides an optimal trade-off.
Number prediction head performance. The number prediction head achieves an overall number prediction accuracy of 53.5%. During the bounding box prediction stage, this accuracy is observed to be 22.84%. This performance gap demonstrates the effectiveness of number prediction head. Besides, the number prediction achieves a mean absolute error (MAE) of 0.51, suggesting high precision as predictions closely align with ground truth values.
Ablations of RDAgent. Table 7 analyzes the effectiveness of the proposed RDAgent. The GPT4-o baseline is established using only Step 3 of RDAgent. We also compare against ReCLIP [36], a two-stage instance ranking method using CLIP [33] for REC tasks. All experiments utilize Faster-RCNN [34] for object detection. Results show that RDAgent consistently outperforms GPT4-o and ReCLIP across all metrics. However, the performance is partially limited by the Faster-RCNN object detector, which achieves only 18.0 mAP on the VisDrone2019-DET [49] dataset.

### v2

@@ 3.4 Dataset comparison
Table 1 compares RefDrone with existing REC datasets. RefDrone stands out for its higher average number of referred objects and use of LMM for expression generation. Its expressions offer richer contextual details than template-based or human-annotated expressions [59]. While RIS-CQ [19] uses LLM for expression generation, it lacks visual content, often producing linguistically complex but visually disconnected expressions. Similarly, RSVG [56] focuses on small targets but lacks expression quality. In contrast, RefDrone comprehensively presents three challenges, establishing it as a challenging benchmark in REC tasks.
@@ 3.5 Evaluation metrics
We introduce instance-level metrics extending traditional REC metrics to address multi-target challenges. Previous benchmarks primarily focus on image-level metrics, which are sufficient for single-target or few-target samples. However, these metrics fail to handle expressions referring to multiple target objects—potentially scaling up to hundreds of instances per expression in our task. To address this limitation, these new metrics provide a more granular evaluation of a model’s capability to identify individual target objects.
Instance-level metrics: Accinst. and F1inst. provide a fine-grained evaluation of the ability to precisely localize individual target objects. We compute the intersection over union (IoU) between the true bounding boxes and the predicted bounding boxes. An IoU $\geq$ 0.5 is considered a true positive (TP), otherwise it is a false positive (FP). Unmatched true bounding boxes are counted as false negatives (FN). For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP). We calculate:
$\text{Acc}_{inst.}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}},$ (1)
@@
Image-level metrics: Accimg. and F1img. evaluate the overall accuracy at the image level. An expression is a true positive (TP) only if all predictions exactly match the true bounding boxes. Any partial or extra prediction is a false positive (FP). For no-target samples, no prediction is counted as a true negative (TN), otherwise a false positive (FP). Image-level metrics use the same formulations as instance-level metrics but are applied to whole expressions rather than individual objects.
4 NGDINO
We introduce Number GroundingDINO (NGDINO) to address multi-target and no-target challenges in REC. Our key insight is that explicit numerical reasoning about target counts enhances model performance on these challenges. NGDINO builds upon GDINO [34] while introducing three components: (1) a number prediction head, (2) learnable number-queries with number-guided query selection, and (3) a number cross-attention module.
NGDINO adopts GDINO’s dual-encoder-single-decoder architecture. The image encoder extracts visual features from input images, while the text encoder processes referring expressions. A feature enhancer facilitates cross-modal fusion, and a language-guided query selection module initializes detection queries. The decoder iteratively refines detection queries through self-attention, image cross-attention, and text cross-attention mechanisms, ultimately producing bounding box predictions. We mainly improve the decoder part, as highlighted by the yellow box in Figure 6.
Number prediction head. We introduce a classification module to predict the number of referred objects $N_{pred}$. This prediction head consists of an FFN layer applied to the detection queries $Q_{det}$, followed by mean pooling:
$N_{prob}=\text{softmax}~{}(~{}\text{MeanPool~{}(~{}FFN}~{}(Q_{det}))),$ (3)
$N_{pred}=\text{argmax}~{}(N_{prob}),$ (4)
where $Q_{det}\in\mathbb{R}^{B\times L_{d}\times D}$ with batch size $B$, detection query length $L_{d}$, and feature dimension $D$. Rather than directly predicting arbitrary counts, we quantize the prediction space into five categories $\mathcal{C}=\{0,1,2,3,4+\}$, where $4+$ represents four or more targets. This quantization addresses the long-tailed distribution of target counts in real-world data (which follows Zipfian distribution [2]).
@@ 4 NGDINO
Number-queries and number-guided query selection. We introduce a learnable set of number-queries $Q_{num}\in\mathbb{R}^{B\times L_{n}\times D}$, where $L_{n}$ is the length of number-queries. $Q_{num}$ is randomly initialized and trained to capture numerical patterns. Based on the predicted target number $N_{pred}$, we select a subset of number-queries $Q_{num}^{sel}$:
$Q_{num}^{sel}=Q_{num}\big{[}:,\ L_{s}\cdot N_{pred}\ :\ L_{s}\cdot(N_{pred}+1),\ :\ \big{]},$ (5)
where $L_{s}$ is the length of selected number-queries. This creates a direct mapping between the predicted number of objects and the specialized queries that encode patterns relevant to that cardinality.
Number cross-attention module. We integrate the number information into the object detection process through a number cross-attention mechanism that operates in parallel with the self-attention in the decoder. In number cross-attention, detection queries $Q_{det}$ serve as queries (Q), while the selected number-queries $Q_{num}^{sel}$ serve as both keys (K) and values (V). The output of this module is added to the self-attention features before subsequent processing in the decoder layer.
Training objective. We adopt GDINO’s bounding box supervision and add a Cross-Entropy loss for number prediction. Ablation studies determine the optimal parameters: selected number-query length $L_{s}=10$ and total number-query length $L_{n}=50$, matching our 5-category quantization scheme.
@@ 5 Experiments
We establish the benchmark with 13 representative models that can perform REC tasks, comprising 3 specialist models and 10 LMMs. The specialist models include MDETR [20], GLIP [25], and GDINO [34]. For the LMMs, we evaluate Shikra [7], ONE-PEACE [47], SPHINX-v2 [27], MiniGPT-v2 [5], Ferret [53], Kosmos-2 [38], Griffon [57], Qwen-VL [3], CogVLM [50], and LLaVA [30]. Detailed model specifications and implementation details are provided in the Appendix.
Table 2: Experimental results of fine-tuning baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks. GDINO-T and GDINO-B denote GroundingDINO [34] with Swin-Tiny and Swin-Base backbones. The best results in each group are denoted with bold.
Methods F1inst. Accinst. F1img. Accimg. MDETR [20] 32.60 19.49 19.17 10.81 GLIP [25] 24.23 14.86 16.92 13.29 GDINO-T [34] 30.50 19.02 29.65 21.07 NGDINO-T (Ours) 33.34 20.98 32.45 22.84 GDINO-B [34] 31.96 19.99 31.69 22.26 NGDINO-B (Ours) 34.44 21.76 34.01 23.89 MiniGPT-v2 [5] 4.97 2.74 13.56 8.97 LLaVA-v1.5 [30] 6.00 3.63 14.43 11.57 Qwen-VL [3] 14.14 7.61 20.10 11.17
Table 3: Experimental results of zero-shot baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks.
@@ 5.1 Experimental results
Fine-tuning results. Table 3 illustrates fine-tuning performance across multiple methods on our RefDrone dataset. The consistently poor performance across all approaches highlights the inherent challenges of the proposed dataset. The specialist model MDETR [20] shows strong performance in instance-level metrics, achieving comparable results to GDINO-B [34], but struggles with image-level understanding. Conversely, LMMs like Qwen-VL [3] achieve superior image-level comprehension while struggling with instance-level tasks. This performance disparity can be attributed to LMMs’ effective global image understanding capabilities but limited effectiveness in detecting small objects due to input resolution constraints (detailed results by object scale are provided in Appendix Table 9).
Our proposed NGDINO consistently outperforms the baseline GDINO [34] across both backbone architectures. Specifically, NGDINO-B achieves state-of-the-art performance, surpassing GDINO-B by significant margins in instance-level metrics (2.48% improvement in F1inst. and 1.77% in Accinst.) as well as image-level metrics (2.32% in F1img. and 1.63% in Accimg.).
Zero-shot results. Table 3 presents the zero-shot evaluation results, assessing the models’ domain generalization capability. CogVLM [50] demonstrates state-of-the-art performance across metrics. However, several advanced models, including Shikra [7], ONE-PEACE [47], and SPHINX-v2 [27], are limited to outputting only a single bounding box, due to constraints in their pre-training data or output strategy. This restriction significantly impacts their performance in multi-target scenarios.
Additional benchmark evaluations. As presented in Table 5, we evaluate our method on the gRefCOCO [28] benchmark with multi-target and no-target samples. The metric Pr@0.5 measures the percentage of predictions with $F1$ = 1 at IoU $\geq$ 0.5, while N-acc denotes the no-target accuracy. NGDINO-T demonstrates a notable improvement over GDINO-T, especially in the N-acc metric, achieving gains of 4.15% on test A and 1.39% on test B, which assess performance on no-target samples. In contrast, the performance gains in Pr@0.5 (0.36% and 0.74% on test A and B, respectively) are more modest. This can be partially attributed to the relatively simple structure of the multi-target queries in gRefCOCO, where expressions predominantly refer to only one or two objects. MDETR [20] achieves higher Pr@0.5 on the test A set. However, this is likely because MDETR tends to produce more outputs, resulting in lower N-acc performance due to an increased number of false positives on no-target examples.
Table 5 demonstrates our method’s effectiveness on the aerial view RSVG [56] benchmark with multi-target samples, where NGDINO-T consistently outperforms previous SOTA methods.
@@
Methods testA testB Pr@0.5 $\uparrow$ N-acc. $\uparrow$ Pr@0.5 $\uparrow$ N-acc. $\uparrow$ MDETR* [20] 50.0 34.5 36.5 31.0 UNINEXT* [52] 46.4 49.3 42.9 48.2 GDINO-T [34] 45.69 79.02 44.83 76.69 NGDINO-T 46.05 83.17 45.57 78.08
Table 5: Experimental results on RSVG [56] dataset. Asterisk (*) denotes results reported in the EarthGPT [58] paper.
Methods Pr@0.5 $\uparrow$ Pr@0.6 $\uparrow$ Pr@0.7 $\uparrow$ Pr@0.8 $\uparrow$ TransVG* [11] 72.41 67.38 60.05 49.10 MGVLF* [56] 76.78 72.68 66.74 56.42 EarthGPT* [58] 76.65 71.93 66.52 56.53 GDINO-T [34] 76.99 75.81 73.36 66.77 NGDINO-T 77.16 76.03 73.81 67.84
Table 6: Ablation study of NGDINO components on the RefDrone dataset with Swin-Tiny backbone.
Number Prediction Head Number Cross-Attention F1inst. Accinst. F1img. Accimg. FPS 30.50 19.02 29.65 21.07 13.5 30.62 19.09 29.90 21.18 12.8 31.25 19.51 30.73 21.71 12.9 33.34 20.98 32.45 22.84 12.3
Table 7: Ablation study on the impact of varying selected number query length. Params indicates additional parameters introduced.
@@
Table 8: Experimental results of two-stage instance ranking methods on the RefDrone benchmark.
Methods F1inst. Accinst. F1img. Accimg. ReCLIP [44] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAnnotator 58.14 41.13 37.07 23.54
@@ 5.2 Ablation studies
Ablation study on NGDINO components. Table 6 presents the analysis of each component in NGDINO. The number prediction head alone shows a slight increase in Accinst. from 19.02% to 19.09%, indicating the minimal impact of this auxiliary head on REC tasks. When introducing the number cross-attention without number selection, we observe a more substantial improvement, with Accinst. increasing to 19.51%. This improvement can be partially attributed to the additional parameters introduced in the decoder. The most significant improvement is achieved through the combination of both the number prediction task and the number cross-attention components. This contribution increases Accinst. to 20.98%. While the additional computational cost results in a minor decrease in inference speed (FPS), this trade-off yields a 1.96% increase in Accinst..
Ablation study on query length. Table 8 analyzes the impact of varying the selected number query length. Utilizing a minimal query length of 1 lacks the capacity to capture complex numerical information. Conversely, extending the query length to 100 increases parameter count and computational overhead, potentially leading to optimization challenges. Through these experiments, we determine that a query length of 10 provides an optimal trade-off.
Performance on number prediction. The number prediction head achieves an overall number prediction accuracy of 53.5%. During the bounding box prediction stage, this accuracy is observed to be 22.84%. This performance gap demonstrates the effectiveness of the number prediction head. Besides, the number prediction achieves a mean absolute error (MAE) of 0.51, suggesting high precision as predictions closely align with ground truth values.
Performance of two-stage instance ranking method. Table 8 analyzes our annotation framework RDAnnotator (without human verification) against alternative two-stage instance ranking methods for REC tasks. We compare: (1) GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ using only Step 3 of RDAnnotator, and (2) ReCLIP [44], which employs CLIP [40] for instance ranking. All experiments utilize Faster-RCNN [41] as the object detector (18.0 mAP on VisDrone2019-DET [60]). Results show RDAnnotator’s ability to produce high-quality annotations. Furthermore, superior instance-level metrics over end-to-end methods reveal the promising potential of two-stage ranking approaches in multi-target challenges.
Figure 7: Some failure cases of NGDINO on RefDrone dataset. Red, green, and yellow boxes indicate true positives, false positives, and false negatives, respectively.